Fermi National Laboratory
 

Accelerator Update for May 11, 2007 - May 14, 2007

 

Friday May 11

The day shift began with the Tevatron (TeV) in a shot setup, with Operations monitoring the Recycler stash (198.1E10), with the Antiproton Source (Pbar) not stacking due to work on a power supply shunt (A:QS603), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

At 8:45 AM, the NuMI beam permit tripped off due to a loss of argon gas flow to beam monitors. Operations contacted experts who had the problem fixed an hour later.

Operations reported at 8:54 AM that the Pbar work ended. There was never a shunt SCR problem; it was a controls problem. Operators put Pbar into stacking mode and established beam to the Accumulator at 9:34 AM.

Operations established store 5421 at 10:33 AM with an initial luminosity of 136.0E30.

At 10:05 AM, the P150 and A150 transfer lines, Booster, and MI critical device controllers (CDC) tripped off. Operators couldn’t find a good reason for this; also a TeV sector D0 utility room detector had tripped off for no apparent reason; everything reset okay.

At 10:15 AM, Operations tried to establish slow spill to the Switchyard (SY) and the Meson MTest experiment, but a septa vacuum problem halted this. Operators and a MI expert investigated.

At 10:32 AM, Operations halted the MI ramp and opened their VCBs to complete controls work on the MI-40 PLC.

Booster RF experts brought an RF station (BRF18) online at 10:39 AM and turned off BRF19.

At 11:23 AM, Operations reported that a QXR problem was holding off slow spill beam.

At 11:36 AM, a Linac RF station (LRF3) tripped off. It reset.

Operations reported at 12:14 PM that they had successfully rebooted QXR and it began to learn in. A Meson expert began tuning the SY beam.

At 12:17 PM, Operators investigated an odor of burnt electronics and found a bad fan on a Booster quadrupole power supply, but the problem was really with a sextupole power supply (SEXL7). Experts began replacing the power supply.

At 12:55 PM, LRF3 tripped off. It reset.

QXR learned in and Operations established beam to MTest around 2:20 PM.

Power supply experts reported that the Booster SEXL7 was back online at 2:25 PM.

At 3:01 PM, LRF3 tripped off for the first of three times in 15 minutes. The first two trips reset, but after the last trip, Linac experts investigated. They had the station back online at 3:55 PM.

Operations reported at 4:16 PM that the Recycler had stashed antiprotons once during the day shift.

At 4:20 PM, LRF3 tripped off. It reset.

MTest reported at 6:25 PM that they were done with beam for the day.

Operations reported that TeV experts began a destructive end-of-store study at 8:30 PM.

Operations terminated store 5421 at 10:31 PM.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Saturday May 12

The midnight shift began with Operations monitoring the stash (137.7E10), with Pbar stacking (37.8E10), and with NuMI and MiniBooNE taking beam.

At 12:11 AM, Operations reported that BRF6 had tripped off many times. It reset.

TeV studies ended at 12:39 AM. Operations prepared D0 for an access.

At 1:03 AM, BRF6 tripped off for the first of five times in about five hours. It reset.

D0 came out of access at 1:32 AM.

Operations put the TeV into shot setup at 2:23 AM.

Operations established store 5424 at 4:57 AM with an initial luminosity of 150.9E30.

At 5:38 AM, Operations reported that BRF6 had begun tripping off so often that operators had bypassed the station and brought BRF19 online in its place.

At 5:48 AM, MIRF17 tripped off and caused NuMI to trip off due to bad beam positions. Everything reset.

At 7:12 AM, Operations reported that Booster tripped off due to a CAMAC crate power supply failure. Operators had the power supply replaced by 7:29 AM.

Operations reported at 8:15 AM that the Recycler had stashed two times during the midnight shift.

At 10:44 AM, LRF2 tripped off. It reset.

At 11:28 AM, BRF2 tripped off. It reset.

At 2:20 PM, Operations received an alarm indicating a MI-60 pond pump trip. This tripped off the MIRF system. The Duty Electrician investigated and reported that the pond pump hadn’t tripped off. Operations thought that a pressure drop had caused the trip. Everything was operating okay.

Operations had the MIRF system back online by 3:07 PM.

MTest shut down for the evening at 7:11 PM.

At 7:22 PM, Operations reported that two Recycler transfers had been lost during the last stash. Operators investigated and discovered a problem with the MI-20 power supplies. They were turned off and the MI-40 power supplies were turned on.

Operations reported at 7:58 PM that the Recycler had stashed four times during the last 12 hours. Accelerator status: store 5424 was at 41.71E30, the stack was at 11.5E10, and the stash was at 176.6E10.

Sunday May 13

The midnight shift began with Operations monitoring the store 5424 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations terminated store 5424 at 2 AM.

Operations put the TeV into shot setup at 2:12 AM.

At 3:13 AM, Operations reported a TeV power supply (T:SF) tripped off due to an overvoltage and fuse fault. Operations called in an expert who reset the fault.

Operations established beam to MTest at 4:18 AM.

At 5:18 AM, LRF3 tripped off. It reset.

At 6:09 AM, KRF2 tripped off. It reset.

Operations put the TeV through a dry squeeze at 6:21 AM.

Operations put the TeV into shot setup at 6:51 AM.

At 7:10 AM, a Booster vacuum pump (IPRF23) tripped off and caused BRF13 & 14 to trip off. Everything reset.

At 7:19 AM, the Booster ORBUMP power supply tripped off. Operators investigated and then called in an expert.

Operations reported at 8:17 AM that the Recycler had stashed three times during the last 12 hours. Accelerator status: the stack was at 45.8E10, and the stash was at 261.6E10.

A Booster expert reported that supply #3 for the ORBUMP had failed. He brought supply #1 online. Booster resumed sending beam at 8:20 AM.

At 9 AM, Operations reported a failed transfer of protons from the MI to the TeV due to a MI beam permit trip. Operators believed that the trip was due to a kicker problem, but before operators could investigate the problem further everything cleared up.

Operations established store 5446 at 10:18 AM with an initial luminosity of 188.27E30.

At 10:22 AM, BRF3 tripped off. It reset.

At 4:13 PM, LRF3 tripped off. It reset.

At 6:15 PM, a Booster vacuum pump (IPRF17) tripped off and caused BRF6 & 7 to trip. Everything reset.

MTest finished using the beam at 7:57 PM.

Operations reported at 8:01 PM that the Recycler had stashed three times during the last 12 hours. Accelerator status: store 5446 was at 71.34E30, the stack was at 31.9E10, and the stash was at 149.9E10.

At 8:32 PM, Operations reported that the deionizer column #2 at CUB alarmed on resistivity. Column #2 closed and column #3 opened up.

Monday May 14

The midnight shift began with Operations monitoring the store 5446 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 2:33 AM, LRF3 tripped off. It reset.

At 4:17 AM, LRF 3 tripped off. It reset.

At 5:55 AM, an air compressor at TeV sector B1 tripped off. Operators reset it.

Operations terminated store 5446 at 8 AM.

Operations reported at 8:01 AM that the Recycler had stashed five times during the last 12 hours. Accelerator status: the stack was at 37.3E10, and the stash was at 304.8E10.

The Plans for Monday

The plans for today are to establish a store this morning and then resume normal stacking, stashing, and storing. The Run Coordinator will allow MI to clean their dirty strainer after the shot and for MI experts to conduct non beam destructive collimator and aperture scans.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/14/2007    email Fermilab